### v2

@@ 3.3 Heterogeneity-Aware Transformer
Following Mixture-of-Size Encoder, the fused representation $\boldsymbol{H}\in\mathbb{R}^{\tilde{T}\times D_{h}}$ is processed by a Heterogeneity-Aware Transformer, a stack of standard Transformer (Vaswani et al., 2017) encoder blocks, yielding representations $\boldsymbol{H}^{\text{en}}\in\mathbb{R}^{\tilde{T}\times D_{h}}$ that are subsequently fed to the prediction head. To inject temporal information into the self-attention mechanism, we propose Dynamic Rotary Position Embedding (DRoPE). Unlike static encodings such as Vaswani et al. (2017); Su et al. (2024), DRoPE adaptively calibrates temporal structure to account for instance-specific periodicities and the varying physical durations of dynamic patches.
Base Structure of DRoPE. Our dynamic position encoding is based on the standard Rotary Position Embedding (RoPE) Su et al. (2024), where a hidden vector $\boldsymbol{z}\in\mathbb{R}^{D_{h}}$ (a query or key) at token index $t\in[1,\tilde{T}$] is treated as a set of $D_{h}/2$ complex numbers. Each pair is rotated by an angle proportional to the position $t$ and a fixed frequency $\theta_{d}$ as below, where $d\in[0,D_{h}/2-1]$ indexes the dimension:
@@
Here the frequencies are traditionally pre-defined as $\theta_{d}=b^{-2d/D_{h}}$. DRoPE generalizes this formulation by dynamically transforming both the frequency $\theta$ and the token position $t$. A detailed derivation of the RoPE calculation is provided in Appendix D.4.
Instance-Level Frequency Modulation. Standard RoPE uses a uniform frequency $\theta$ for all sequences, failing to capture the unique periodic structures of heterogeneous time series (analysis in Theorem D.1). We propose to modulate the base frequencies using instance-specific spectral features. Specifically, given a time series input $\mathbf{X}$, we extract the magnitudes of the first $\omega$ low-frequency components via Fast Fourier Transform, denoted as $\mathbf{X}_{\mathrm{fft}}$. A lightweight MLP maps these features to a scaling factor $\gamma$ and a bias $\beta$, to produce adaptive frequencies $\theta_{\mathrm{inst}}$:
$\theta_{\text{inst},d}=\gamma_{d}\cdot\theta_{d}+\beta_{d}\penalty 10000\ ,\text{where\penalty 10000\ }\gamma,\beta=\mathrm{MLP}(\mathbf{X}_{\mathrm{fft}})$ (5)
This affine transformation allows Kairos to re-scale the temporal rotation based on the dominant information like trends and seasonalities in the specific time series instance.
Granularity-Aware Position Calibration. In our multi-granularity framework, the discrete token index $t$ is no longer an accurate proxy for time because different tokens represent different physical durations, a direct consequence of the granularity routing and tokenization in our Mixture-of-Size Encoder, as described in Section 3.2. To maintain temporal consistency across mixed patch sizes, we replace the integer index $t$ with a calibrated position $\tilde{t}=\sum_{k=1}^{t-1}{\bar{P}_{k}}/{\bar{P}_{\mathrm{g\_min}}}$, where $\bar{P}_{k}$ is the patch size of the $k$-th token and $\bar{P}_{\mathrm{g\_min}}=\min\{P_{i}\}_{i=1}^{G}$ is the globally minimum patch size of the finest granularity.
The resulting DRoPE operation unifies these two components, rotating the hidden vector $\boldsymbol{z}$ at position $t$ as follows:
@@ 3.4 Multi-Patch Decoder
Each embedding $\boldsymbol{h}^{\text{de}}_{j}$ is then projected into the observation space via a shared prediction head, which is implemented as a residual FFN block Das et al. (2024), to generate a future patch of length $H_{s}=\lceil H/J\rceil$ as
@@
This multi-token design is strategically for two objectives regarding predominant forecasting paradigms. (i) Error mitigation and efficiency: Compared to conventional single-step autoregressive models (Ansari et al., 2024; Liu et al., 2024c), predicting $J$ patches in parallel improves computational efficiency and mitigates cumulative error by reducing iterative decoding steps. (ii) Horizon flexibility: Unlike models that generate a single larger patch (Das et al., 2024), our approach offers superior granularity for variable prediction horizons. By assembling the forecast from the specific subset of tokens required, Kairos ensures high-fidelity alignment with the target length and avoids the imprecise truncation associated with oversized outputs. We provide a detailed analysis of this design in Appendix B. By combining the Mixture-of-Size Encoder for adaptive input processing with this Multi-Patch Decoder for flexible output generation, Kairos achieves a unified and parameter-efficient framework for zero-shot time series forecasting.
@@ 3.5 Training Corpus
For pre-training, we curated the Predictability-Stratified Time Series (PreSTS) corpus, comprising over 300 billion time points. PreSTS integrates large-scale real-world time series from diverse domains and a complementary synthetic dataset to ensure broad coverage. All test sets from the GIFT-Eval (Aksu et al., 2024) and TSLib (Wang et al., 2024b) benchmarks are explicitly excluded to prevent data leakage. While dataset diversity is widely regarded as essential for TSFM training, recent studies indicate that low-predictability or anomalous sequences can substantially impair forecasting (Cheng et al., 2024). To address this, we stratify real-world datasets into five tiers based on predictability and assign higher sampling probabilities to more predictable data during training. This strategy promotes stable and high-quality pre-training without compromising data diversity. Further details are provided in Appendix E.1.
@@ 4 Experiment
Table 1: Performance evaluation on the GIFT-Eval. We evaluate using normalized MASE and CRPS (Section 4.1.1). Baselines encompass statistical, Deep Learning (DL), and TSFM methods, with TSFMs categorized into TestData Leakage and Zero-Shot following official protocols. TestData Leakage signifies that the model’s training set included test data, whereas Zero-Shot indicates that the model was trained without any data from the test set or its corresponding training split. Baseline results are officially reported by GIFT-Eval.
@@
Figure 3: Zero-shot performance on TSLib averaged over prediction lengths {96, 192, 336, 720}. The subscripts $l$, $b$, $s$, $m$, and $a$ represent model sizes of large, base, small, mini and advanced, respectively. The complete experimental results are presented in Appendix F.
@@ 4.1 Experiment Settings
4.1.1 Evaluation Datasets and Metric
To comprehensively evaluate our proposed method, we conducted zero-shot assessments on two well-known public benchmarks: GIFT-Eval (Aksu et al., 2024) and Time-Series-Library (TSLib) (Wang et al., 2024b). We train Kairos in three sizes: mini (10M), small (23M) and base (53M). The training details are provided in Appendix D.1.
Benchmarks. The GIFT-Eval benchmark comprises 97 tasks spanning short- (55), medium- (21), and long-term (21) horizons, facilitating a comprehensive assessment. For TSLib, we selected the ETT and Weather (Zhou et al., 2021) datasets, consistent with Time-MoE (Shi et al., 2024). To test adaptability to diverse frequencies, we also included the Saugeen datasets (Godahewa et al., 2021), denoted as Saugeen (D) and Saugeen (W) for daily and weekly intervals, respectively. The detailed description of the evaluation datasets, splitting configurations, and evaluation length selection are provided in Appendix E.2 and E.3.
Metric. In our evaluation of GIFT-Eval, consistent with prior research (Auer et al., 2025; Liu et al., 2025), we employ the official Mean Absolute Scaled Error (MASE) and Continuous Ranked Probability Score (CRPS) metrics to assess point and probabilistic forecasting performance, respectively. These metrics are first normalized by the Seasonal Naïve baseline for each individual task. The final score is then computed as the geometric mean of these normalized values. For the TSLib benchmark, we followed previous works (Liu et al., 2023; 2024c; Nie et al., 2023) by adopting Mean Squared Error (MSE) and Mean Absolute Error (MAE) to evaluate time series forecasting performance. The metric calculations are detailed in Appendix D.2.1.
4.1.2 Compared Methods
We compare Kairos with several state-of-the-art models, including TSFMs such as Sundial (Liu et al., 2025), Toto (Cohen et al., 2025), YingLong (Wang et al., 2025), VisionTS (Chen et al., 2025), Time-MoE (Shi et al., 2024), ChronosBolt (Ansari et al., 2024), TTM (Ekambaram et al., 2024), Moirai (Woo et al., 2024), Timer-XL (Liu et al., 2024c), TimesFM-2.0 (Das et al., 2024), Chronos (Ansari et al., 2024), and advanced full-shot deep learning models, including PatchTST (Nie et al., 2023), DLinear (Zeng et al., 2022), iTransformer (Liu et al., 2023), Pathformer (Chen et al., 2024), and TimesNet (Wu et al., 2022).
@@ 4.2 Zero-shot Evaluation
Finding 1: Kairos achieves superior zero-shot forecasting performance with significantly higher parameter efficiency. On the GIFT-Eval benchmark, as presented in Table 1, $\textsc{Kairos}_{\text{base}}$ (53M) achieves the best MASE and the second-best CRPS among state-of-the-art task-specific deep learning methods and other TSFMs. Notably, $\textsc{Kairos}_{\text{small}}$ (23M) surpasses both Toto and Sundial in MASE, despite having a parameter count that is 6.6 and 5.6 times smaller, respectively. On the TSLib benchmark, as shown in Figure 3, our lightweight $\textsc{Kairos}_{\text{mini}}$ (10M) surpasses the performance of both recent advanced TSFMs and the majority of full-shot deep learning models. These results confirm that Kairos delivers a significant improvement in predictive capability without relying on massive scale.
@@ 4.3 Ablation Study
Finding 2: The synergistic integration of Mixture-of-Size Encoding, DRoPE, and the Multi-Patch Decoding is critical for adapting to heterogeneous time series and achieving superior forecasting performance. Table 2 details the component-wise contributions. (i) Encoder Design: Kairos outperforms variants using fixed patch sizes or excluding null experts, confirming that dynamic, sparse multi-granularity modeling is critical for capturing diverse temporal patterns. (ii) Positional Embedding: DRoPE proves superior to standard RoPE (Su et al., 2024). Our ablation further isolates the gains from Instance-Level Frequency Modulation and Granularity-Aware Position Calibration, verifying the benefit of adapting to both spectral heterogeneity and physical time distortion caused by dynamic patching. (iii) Decoding Strategy: The Multi-Patch Decoding shows clear superiority over standard single-patch autoregressive decoding. The integration of both components yielded the best performance across 97 GIFT-Eval tasks, which validates the superiority of our designs. Furthermore, the performance gap widens as the forecast horizon increases, demonstrating Kairos’s superior long-term forecasting capability.
Table 2: Ablation study comparing encoder design, positional embedding and decoder strategy. Models were evaluated using the normalized MASE (detailed in Section 4.1.1) across individual prediction horizons and as an aggregate across all tasks.
@@ 4.4.1 Mixture-of-Size Tokenization Analysis
Finding 3: Kairos demonstrates adaptive capability to varying information densities, enabling it to model complex temporal dynamics with appropriate granularity. Instead of enforcing a uniform resolution, Kairos dynamically routes segments to optimal granularity based on local information. As visualized in Figure 4, Kairos strategically allocates fine-grained tokens to regions with abrupt changes or high volatility, while employing coarse-grained tokens for stable trends. This adaptive allocation mechanism reduces computational redundancy on low-information segments, allowing the model to achieve superior performance with a compact parameter budget compared to rigid architectures. See Appendix C for a detailed inference speed analysis.
@@ 4.4.2 DRoPE Analysis
Finding 4: Modulating RoPE frequencies $\theta$ instance-wise can indeed better model time series temporal relationships. To verify that DRoPE’s performance gains stem from its instance-specific $\theta_{\text{inst}}$ modulation, we designed control experiments that disrupt this mechanism (Table 3; Appendix E.4): (i) Intra-Dataset Shuffle: randomly permute $\theta_{\text{inst}}$ across instances within the same dataset; (ii) Inter-Dataset Shuffle: assign $\theta_{\text{inst}}$ from instances of different datasets; (iii) Fixed RoPE: use standard RoPE without any modulation. The results reveal a clear performance gradient (DRoPE $\succ$ Intra-Shuffle $\succ$ Fixed RoPE $\succ$ Inter-Shuffle), indicating that DRoPE learns instance-specific representations to modulate $\theta$. Intra-Dataset Shuffle outperforms Fixed RoPE, as samples within the same dataset share similar spectral characteristics, allowing shuffled $\theta_{\text{inst}}$ to remain partially compatible. However, Intra-Dataset Shuffle still incurs a 1.76% degradation compared to DRoPE, suggesting that instance-level adaptation contributes to the performance gains beyond dataset-level statistics. Inter-Dataset Shuffle yields the worst results (28.32% degradation), as $\theta_{\text{inst}}$ from different datasets cannot capture the target series’ temporal dynamics. These results collectively validate the effectiveness of instance-level $\theta_{\text{inst}}$ adaptation in DRoPE.
@@ 5 Conclusions
We introduce Kairos, a parameter-efficient foundation model explicitly tailored for time series heterogeneity. Its architecture features three key innovations: a Mixture-of-Size Encoder for adaptive granularity, a Heterogeneity-Aware Transformer with Dynamic RoPE for instance-specific temporal calibration, and a Multi-Patch Decoder for efficient parallel forecasting. Extensive experiments demonstrate that Kairos achieves superior performance, validating the necessity of aligning architectural inductive biases with the structural nuances of time series data. One limitation of our current approach is that Kairos relies on channel-independent modeling (Nie et al., 2023) to support multivariate time series and does not explicitly capture inter-variable dependencies. Future work aims to model these interactions and extend Kairos to tasks beyond forecasting.
Impact Statement
Appendix A Etymology of Kairos
Kairos was chosen for its mythological significance as the Greek god of the “right or critical moment”, reflecting our model’s ability to select the ideal set of granularities and positional encoding for time series characterized by varying information densities and frequencies. Furthermore, this continues a thematic tradition seen in other foundational models we benchmark against, such as Chronos (Ansari et al., 2024) (the personification of time) and Moirai (Woo et al., 2024) (the personifications of destiny).
Appendix B Analysis of Multi-Patch Prediction

### v3

@@ 3.3 Heterogeneity-Aware Transformer
Following Mixture-of-Size Encoder, the fused representation $\boldsymbol{H}\in\mathbb{R}^{\tilde{T}\times D_{h}}$ is processed by a Heterogeneity-Aware Transformer, a stack of standard Transformer [36] encoder blocks, yielding representations $\boldsymbol{H}^{\text{en}}\in\mathbb{R}^{\tilde{T}\times D_{h}}$ that are subsequently fed to the prediction head. To inject temporal information into the self-attention mechanism, we propose Dynamic Rotary Position Embedding (DRoPE). Unlike static encodings such as Vaswani et al. [36], Su et al. [35], DRoPE adaptively calibrates temporal structure to account for instance-specific periodicities and the varying physical durations of dynamic patches.
Base Structure of DRoPE. Our dynamic position encoding is based on the standard Rotary Position Embedding (RoPE) [35], where a hidden vector $\boldsymbol{z}\in\mathbb{R}^{D_{h}}$ (a query or key) at token index $t\in[1,\tilde{T}$] is treated as a set of $D_{h}/2$ complex numbers. Each pair is rotated by an angle proportional to the position $t$ and a fixed frequency $\theta_{d}$ as below, where $d\in[0,D_{h}/2-1]$ indexes the dimension:
@@
Here the frequencies are traditionally pre-defined as $\theta_{d}=b^{-2d/D_{h}}$. DRoPE generalizes this formulation by dynamically transforming both the frequency $\theta$ and the token position $t$. A detailed derivation of the RoPE calculation is provided in Appendix H.4.
Instance-Level Frequency Modulation. Standard RoPE uses a uniform frequency $\theta$ for all sequences, failing to capture the unique periodic structures of heterogeneous time series (analysis in Theorem H.1). We propose to modulate the base frequencies using instance-specific spectral features. Specifically, given a time series input $\mathbf{X}$, we extract the magnitudes of the first $\omega$ low-frequency components via Fast Fourier Transform, denoted as $\mathbf{X}_{\mathrm{fft}}$. A lightweight MLP maps these features to a scaling factor $\gamma$ and a bias $\beta$ to produce adaptive frequencies $\theta_{\mathrm{inst}}$. To accommodate the exponentially decaying nature of RoPE’s base frequencies, we perform this affine transformation in log-space:
$\log\theta_{\text{inst},d}=\gamma_{d}\cdot\log\theta_{d}+\beta_{d}\penalty 10000\ ,\text{where\penalty 10000\ }\gamma,\beta=\mathrm{MLP}(\mathbf{X}_{\mathrm{fft}})$ (5)
This log-space adaptation allows Kairos to robustly re-scale the temporal rotation based on the dominant information like trends and seasonalities in the specific time series instance (see Appendix H.4.2 for a detailed discussion).
Granularity-Aware Position Calibration. In our multi-granularity framework, the discrete token index $t$ is no longer an accurate proxy for time because different tokens represent different physical durations, a direct consequence of the granularity routing and tokenization in our Mixture-of-Size Encoder, as described in Section 3.2. To maintain temporal consistency across mixed patch sizes, we replace the index $t$ with a calibrated position $\tilde{t}=\sum_{k=1}^{t-1}{\bar{P}_{k}}/{\bar{P}_{\mathrm{g\_min}}}$, where $\bar{P}_{k}$ is the patch size of the $k$-th token and $\bar{P}_{\mathrm{g\_min}}=\min\{P_{i}\}_{i=1}^{G}$ is the globally minimum patch size of the finest granularity.
Unifying these two components, rotating the hidden vector $\boldsymbol{z}$ at position $t$ as follows:
@@ 3.4 Multi-Patch Decoder
Each embedding $\boldsymbol{h}^{\text{de}}_{j}$ is then projected into the observation space via a shared prediction head, which is implemented as a residual FFN block [11], to generate a future patch of length $H_{s}=\lceil H/J\rceil$ as
@@
This multi-token design serves two primary objectives: (i) error mitigation and computational efficiency via parallel patch prediction, and (ii) horizon flexibility for exact alignment with various target lengths. We provide a detailed comparative analysis of this design in Appendix D. By pairing this flexible Multi-Patch Decoder with the adaptive Mixture-of-Size Encoder, Kairos achieves a unified and parameter-efficient framework for zero-shot time series forecasting.
@@ 3.5 Training Corpus
For pre-training, we curated the Predictability-Stratified Time Series (PreSTS) corpus, comprising over 300 billion time points. PreSTS integrates large-scale real-world time series from diverse domains and a complementary synthetic dataset to ensure broad coverage. All test sets from the GIFT-Eval [3] and TSLib [39] benchmarks are explicitly excluded to prevent data leakage. While dataset diversity is widely regarded as essential for TSFM training, recent studies indicate that low-predictability or anomalous sequences can substantially impair forecasting [9]. To address this, we stratify real-world datasets into five tiers based on predictability and assign higher sampling probabilities to more predictable data during training. This strategy promotes stable and high-quality pre-training without compromising data diversity. Further details on dataset construction, along with an experimental analysis decoupling the contribution of the corpus from our model architecture, are provided in Appendix I.1 and Appendix G, respectively.
@@ 4 Experiment
Table 1: Performance evaluation on the GIFT-Eval. We evaluate using normalized MASE and CRPS (Section 4.1). Baselines encompass statistical, Deep Learning (DL), and TSFM methods, with TSFMs categorized into TestData Leakage and Zero-Shot following official protocols. TestData Leakage signifies that the model’s training set included test data, whereas Zero-Shot indicates that the model was trained without any data from the test set or its corresponding training split. Baseline results are officially reported by GIFT-Eval.
@@
Figure 3: Zero-shot performance on TSLib averaged over prediction lengths {96, 192, 336, 720}. The subscripts $l$, $b$, $s$, $m$, and $a$ represent model sizes of large, base, small, mini and advanced, respectively. The complete experimental results are presented in Appendix J.
@@ 4.1 Experiment Settings
To comprehensively evaluate our proposed method, we conducted zero-shot assessments on two well-known public benchmarks: GIFT-Eval [3] and Time-Series-Library (TSLib) [39]. We train Kairos in three sizes: mini (10M), small (23M) and base (53M). The training details are provided in Appendix H.1.
Benchmarks. The GIFT-Eval benchmark comprises 97 tasks spanning short- (55), medium- (21), and long-term (21) horizons, facilitating a comprehensive assessment. For TSLib, we selected the ETT and Weather [46] datasets, consistent with Time-MoE [34]. To test adaptability to diverse frequencies, we included the Saugeen datasets [15], denoted as Saugeen (D) and Saugeen (W) for daily and weekly intervals, respectively. Details of the compared methods, evaluation datasets, splitting configurations, and evaluation length selection are provided in Appendices I.2, I.3, and I.4.
Metric. In our evaluation of GIFT-Eval, consistent with prior research [5, 28], we employ the official Mean Absolute Scaled Error (MASE) and Continuous Ranked Probability Score (CRPS) metrics to assess point and probabilistic forecasting performance, respectively. These metrics are first normalized by the Seasonal Naïve baseline for each individual task. The final score is then computed as the geometric mean of these normalized values. For the TSLib benchmark, we followed previous works [26, 27, 29] by adopting Mean Squared Error (MSE) and Mean Absolute Error (MAE) to evaluate time series forecasting performance. The metric calculations are detailed in Appendix H.2.1.
@@ 4.2 Zero-shot Evaluation
Finding 1: Kairos achieves superior zero-shot forecasting performance with significantly higher parameter efficiency and learns highly transferable representations.. On the GIFT-Eval benchmark, as presented in Table 1, $\textsc{Kairos}_{\text{base}}$ (53M) achieves the best MASE and the second-best CRPS among state-of-the-art task-specific deep learning methods and other TSFMs. Notably, $\textsc{Kairos}_{\text{small}}$ (23M) surpasses both Toto and Sundial in MASE, despite having a parameter count that is 6.6 and 5.6 times smaller, respectively. On the TSLib benchmark, as shown in Figure 3, our lightweight $\textsc{Kairos}_{\text{mini}}$ (10M) surpasses the performance of both recent advanced TSFMs and the majority of full-shot deep learning models. These results confirm that Kairos delivers a significant improvement in predictive capability without relying on massive scale. Furthermore, we demonstrate in Appendix C that these learned representations are highly transferable to downstream classification tasks.
@@ 4.3 Ablation Study
Table 2: Ablation study comparing encoder design, positional embedding and decoder strategy. Models were evaluated using the normalized MASE (detailed in Section 4.1) across prediction horizons and as an aggregate across all tasks.
@@
Finding 2: The synergistic integration of Mixture-of-Size Encoding, DRoPE, and the Multi-Patch Decoding is critical for adapting to heterogeneous time series and achieving superior forecasting performance. Table 2 details the component-wise contributions. (i) Encoder Design: Kairos outperforms variants using fixed patch sizes or excluding null experts, confirming that dynamic, sparse multi-granularity modeling is critical for capturing diverse temporal patterns. (ii) Positional Embedding: DRoPE proves superior to standard RoPE [35]. Our ablation further isolates the gains from Instance-Level Frequency Modulation and Granularity-Aware Position Calibration, verifying the benefit of adapting to both spectral heterogeneity and physical time distortion caused by dynamic patching. (iii) Decoding Strategy: The Multi-Patch Decoding shows clear superiority over standard single-patch autoregressive decoding. The integration yielded the best performance across 97 GIFT-Eval tasks, which validates the superiority of our designs. Furthermore, the performance gap widens as the forecast horizon increases, demonstrating Kairos’s superior long-term forecasting capability.
@@ 4.4.1 Mixture-of-Size Tokenization Analysis
Finding 3: Kairos demonstrates adaptive capability to varying information densities, enabling it to model complex temporal dynamics with appropriate granularity. Instead of enforcing a uniform resolution, Kairos dynamically routes segments to optimal granularity based on local information. As visualized in Figure 4, Kairos strategically allocates fine-grained tokens to regions with abrupt changes or high volatility, while employing coarse-grained tokens for stable trends. This adaptive allocation mechanism reduces computational redundancy on low-information segments, allowing the model to achieve superior performance with a compact parameter budget compared to rigid architectures. See Appendix F for a detailed inference speed analysis. Furthermore, to demonstrate that Kairos learns genuine segment-level adaptation rather than merely relying on a generic local multi-scale inductive bias, we provide additional experimental evidence in Appendix E.
@@ 4.4.2 DRoPE Analysis
Finding 4: Modulating RoPE frequencies $\theta$ instance-wise can indeed better model time series temporal relationships. To verify that DRoPE’s performance gains stem from its instance-specific $\theta_{\text{inst}}$ modulation, we designed control experiments that disrupt this mechanism (Table 3; Appendix I.5): (i) Intra-Dataset Shuffle: randomly permute $\theta_{\text{inst}}$ across instances within the same dataset; (ii) Inter-Dataset Shuffle: assign $\theta_{\text{inst}}$ from instances of different datasets; (iii) Fixed RoPE: use standard RoPE without any modulation. The results reveal a clear performance gradient (DRoPE $\succ$ Intra-Shuffle $\succ$ Fixed RoPE $\succ$ Inter-Shuffle), indicating that DRoPE learns instance-specific representations to modulate $\theta$. Intra-Dataset Shuffle outperforms Fixed RoPE, as samples within the same dataset share similar spectral characteristics, allowing shuffled $\theta_{\text{inst}}$ to remain partially compatible. However, Intra-Dataset Shuffle still incurs a 1.76% degradation compared to DRoPE, suggesting that instance-level adaptation contributes to the performance gains beyond dataset-level statistics. Inter-Dataset Shuffle yields the worst results (28.32% degradation), as $\theta_{\text{inst}}$ from different datasets cannot capture the temporal dynamics of the target series. These results collectively validate the effectiveness of instance-level $\theta_{\text{inst}}$ adaptation in DRoPE.
@@ 5 Conclusions
We introduce Kairos, a parameter-efficient foundation model explicitly tailored for time series heterogeneity. Its architecture features three key innovations: a Mixture-of-Size Encoder for adaptive granularity, a Heterogeneity-Aware Transformer with Dynamic RoPE for instance-specific temporal calibration, and a Multi-Patch Decoder for efficient parallel forecasting. Extensive experiments demonstrate that Kairos achieves superior performance, validating the necessity of aligning architectural inductive biases with the structural nuances of time series data. One limitation of our current approach is that Kairos relies on channel-independent modeling [29] to support multivariate time series and does not explicitly capture inter-variable dependencies. Future work aims to model these interactions and extend Kairos to tasks beyond forecasting.
Appendix A Limitations and Future Work
While Kairos demonstrates superior parameter efficiency and zero-shot performance, we acknowledge several limitations that provide avenues for future research.
Channel-Independent Modeling. Currently, Kairos adopts a channel-independent modeling approach [29] to handle multivariate time series, treating each variable as an individual sequence without explicitly modeling the relationships between variables. However, our architectural innovations are orthogonal to the channel-independence assumption. It is easy to incorporate channel dependency modeling for further enhancement. For instance, a channel-mixer block can be easily integrated to explicitly capture cross-channel correlations. We plan to explore these extensions in future work.
Evaluation Scope. Our empirical evaluation primarily focuses on zero-shot forecasting, which is the primary evaluation setting for TSFMs [28, 10]. As a generative task, forecasting directly reflects whether a model can capture temporal structures such as trend, seasonality, and local dynamics. Furthermore, we provide evaluation results on classification tasks in Appendix C, which demonstrate that Kairos successfully learns transferable representations. We plan to further extend Kairos to tasks such as anomaly detection and imputation in future versions.
Appendix B Etymology of Kairos
Kairos was chosen for its mythological significance as the Greek god of the “right or critical moment”, reflecting our model’s ability to select the ideal set of granularities and positional encoding for time series characterized by varying information densities and frequencies. Furthermore, this continues a thematic tradition seen in other foundational models we benchmark against, such as Chronos [4] (the personification of time) and Moirai [40] (the personifications of destiny).
Appendix C Transferability on Classification Tasks
To verify whether Kairos learns generalizable features that transfer beyond forecasting tasks, we evaluate its performance on downstream time series classification.
Experimental Setup. Following MOMENT [16], we evaluate on 91 datasets from the UCR archive [12]. To isolate representation quality, Kairos-Base and MOMENT use a frozen encoder with a lightweight classifier, comparing them against fully fine-tuned baselines (GPT4TS [47], TimesNet [41]).
Results. The classification results are summarized in Table 4. Despite relying solely on frozen representations, Kairos consistently outperforms both MOMENT and the fully fine-tuned baselines. These results provide evidence that our architectural innovations do not merely overfit to forecasting dynamics. Instead, by addressing temporal heterogeneity, Kairos extracts high-level semantic features that are highly transferable to adaptation-heavy settings like classification. This solidifies Kairos’s capacity as a general-purpose foundation model capable of supporting diverse downstream time series applications.
Table 4: Representation transfer results on 91 UCR classification datasets. Baseline results follow MOMENT [16].
Model Accuracy $\uparrow$ Rank Mean $\downarrow$ Rank Std $\downarrow$
MOMENT (Frozen) 0.794 1.90 0.80
GPT4TS (Fine-tuned) 0.567 3.27 0.81
TimesNet (Fine-tuned) 0.573 3.32 0.74
[…]
Appendix D Analysis of Multi-Patch Prediction